OpenStudy (anonymous):

Should an estimate or the exact answer be calculated? Then do the appropriate calculation. The conservation club participated in a beach cleanup program over a 5 week period. Their goal was to collect at least 60 pounds of trash. They collected pounds in each of the first two weeks and pounds for each of the last three weeks. Did the club reach its goal? A. An estimate can be used to know if the goal was reached. 12 • 2 + 13 • 3 = 63 At least 63 lb were collected. B. The exact answer is needed to know if the goal was reached. pounds were collected.

OpenStudy (anonymous):

medal

OpenStudy (ineedserioushelp):

Is there something missing here: They collected pounds in each of the first two weeks and pounds for each of the last three

OpenStudy (anonymous):

12 3/4 13 7/8

OpenStudy (ineedserioushelp):

Well then the answer is A. Because we know how many pounds they collected in all the weeks, all we have to do is add it up and we have an estimate of how much they collected. And because we know their goal was 60 pounds, we know that they have not only reached their goal but actually surpassed it.

OpenStudy (anonymous):

can u help on a other 1

OpenStudy (ineedserioushelp):

Sure, I'll try :p

OpenStudy (anonymous):

Should an estimate or the exact answer be calculated? Then do the appropriate calculation. Megan had a $50 gift card for the sporting goods shop. She made 3 purchases of $6.75 each and 2 purchases of $10.50 each. How much money is left on the card? A. An estimate is sufficient. There is about $9.00 left on the card. B. The exact answer is needed. There is $8.75 left on the card.

OpenStudy (ineedserioushelp):

Oh, sorry for taking so long, I didn't notice you posted it yet. Okay, we will need an exact answer here, so the answer is B.

OpenStudy (anonymous):

and one more

OpenStudy (anonymous):

An electrician has a spool of wire 200 m long. She needs four pieces of wire measuring 42.3 m, 37.8 m, 55.6 m, and 35.7 m. How much wire will be left on the spool after she finishes her job? A. 20 m B. 28 m C. 28.6 m D. 29.6 m

OpenStudy (ineedserioushelp):

Sure, why not

OpenStudy (anonymous):

there

OpenStudy (anonymous):

it is posted

OpenStudy (ineedserioushelp):

Okay one sec

OpenStudy (ineedserioushelp):

Well, to figure this one out we need to do some subtraction: 200 - 42.3 = 157.7 , 157.7 - 37.8 = 119.9 , 119.9 - 55.6 = 64.3 , 64.3 - 35.7 = 28.6 Answer is C. 28.6m

OpenStudy (anonymous):

thanx

OpenStudy (ineedserioushelp):

No problem, good luck!